### v1

# Can Quasi Periodic Oscillations Encode Traces of Black Hole Phase Transitions ?

Abstract
Finding observational evidence or imprints of black hole phase transitions is a promising and active area of research. Aiming to contribute to this direction, in this work, we probe the well-known thermodynamic phase structure of the Reissner Nordström Anti de Sitter (RN-AdS) black hole through the lens of its quasi-periodic oscillations (QPOs). Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational patterns? These were the central questions guiding our study. By analyzing the upper and lower QPO frequencies within the Relativistic Precession (RP) model, we observe that the frequency-temperature relationship reflects distinguishing features corresponding to the small, intermediate, and large black hole phases, offering insights into their stability properties. The analysis is further extended to other QPO models, including the Warped Disk (WD) and Epicyclic Resonance (ER) models. The presence of distinct phase structures remains evident across all models which is an encouraging outcome that may be worth exploring further in the context of other black hole systems.
I Introduction
Black holes are among the most enigmatic and intriguing consequences predicted by General Relativity (GR), highlighting the theory’s profound impact on our understanding of gravity. Since the formulation of GR by Einstein in 1915, it has served as a solid theoretical framework for describing the curvature of spacetime. A pivotal confirmation of GR came with the detection of gravitational waves from black hole mergers by the LIGO collaboration, a groundbreaking achievement in the field of gravitational physics [2]. This was followed by the historic imaging of supermassive black holes by the Event Horizon Telescope (EHT), first capturing the black hole in the M87 galaxy, and later in Sagittarius A* (SgrA*) at the center of our own Milky Way [3, 4, 5, 6, 7, 8]. These revolutionary images unveiled a dark central region surrounded by a glowing photon ring, the characteristics of which provide valuable insights into the nature of black holes and the gravitational framework governing them [9, 10, 11, 12].
In the early 1970s, pivotal research in black hole thermodynamics laid the foundation for understanding the connection between black holes and the laws of thermodynamics [13, 14, 15, 16]. This foundational work has since been expanded upon by numerous studies, revealing a wealth of fascinating insights into this field [17, 18, 19, 20, 21, 22, 23]. One intriguing aspect of black hole thermodynamics is the exploration of phase transitions [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Black holes can exhibit various types of phase transitions, including the Davies type phase transition [24], the Hawking–Page phase transition [25], the extremal phase transition (which describes the transition of black holes from non-extremal to extremal states) [26, 27, 28, 29, 30, 31, 32, 33, 34], and the Van der Waals-type phase transition (which is similar to the phase transition seen in Van der Waals fluids) [35, 36, 37, 38, 39, 40, 41].
Quasiperiodic oscillations (QPOs), detected in the X-ray flux from black holes and neutron stars, have emerged as powerful tools for probing the physics of strong gravity regimes. Characterized by brightness fluctuations at nearly periodic intervals, these oscillations are widely believed to originate from fundamental processes within the accretion disk, including relativistic effects and complex disk dynamics. In particular, the observation of twin-peak QPOs in several compact object systems has spurred extensive theoretical and observational studies, often attributing these features to resonances or intrinsic oscillatory behavior within the disk structure. As observational techniques continue to improve, the demand for more sophisticated theoretical frameworks becomes increasingly evident. Since their first identification through spectral and temporal studies in X-ray binaries [42], QPOs have been examined from a range of perspectives. Among various theoretical models, those focusing on the motion of particles in curved spacetime have gained significant attention. In such scenarios, the modulated trajectories of charged test particles are considered central to shaping the accretion flow and driving the observed oscillatory signals [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Recent computational efforts have employed general relativistic hydrodynamic simulations to explore QPO formation mechanisms around black holes, particularly in spacetimes such as Kerr and hairy black holes [58]. These studies suggest that perturbations in the accreting plasma can generate spiral shock patterns, which are intimately connected to the emergence of QPOs [59, 60, 61]. Likewise, models based on Bondi–Hoyle–Lyttleton accretion dynamics show that shock cones, formed due to the gravitational influence of compact objects, can produce distinct QPO signatures [62, 63, 64, 65, 66]. Such approaches have not only succeeded in reproducing the QPO characteristics of well-known sources like GRS 1915+105 [67], but also offer predictive insight into oscillatory behavior near supermassive black holes, including the one in M87 [68]. The dynamics of test particle motion and the resulting QPOs around black holes remain a rich area of investigation key contributions can be found in Refs.[69, 70, 71, 72, 73, 74, 75].
The static, spherically symmetric metric of RN AdS black hole is of the form
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (1)
with the lapse function $f(r)$ given by
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{\kappa^{2}},$ (2)
We set the AdS length $\kappa=1$ for the rest of our analysis. Here $M$ and $Q$ represent the ADM mass and electric charge of the black hole, respectively.
The mass parameter $M$ can be expressed in terms of the event horizon radius $r_{+}$ as
$M=\frac{Q^{2}+r_{+}^{4}+r_{+}^{2}}{2r_{+}}$ (3)
The temperature of the black hole can be calculated by using the formula,
$T=\frac{1}{4\pi}f^{\prime}(r)_{r=r_{+}}=\frac{-Q^{2}+3r_{+}^{4}+r_{+}^{2}}{4\pi r_{+}^{3}}$ (4)
The critical values associated with the system can be found by solving the following equations
$\frac{dT}{dr_{+}}=0\quad,\quad\frac{d^{2}T}{dr_{+}^{2}}=0$ (5)
The solution of the thermodynamic equations reveals the existence of a critical charge given by
$Q_{C}=\frac{1}{\sqrt{6}},$ (6)
below which ($Q<Q_{C}$) the RN-AdS black hole exhibits a Van der Waals (VdW)-like phase transition. In this regime, the system displays three distinct black hole phases: a small black hole (SBH) branch, an intermediate black hole (IBH) branch, and a large black hole (LBH) branch. For $Q>Q_{C}$, the phase transition behavior disappears. The local thermodynamic stability of these black hole phases can be analyzed through the sign of the specific heat, defined by
$C=\frac{dM}{dT},$ (7)
where $M$ is the mass and $T$ is the Hawking temperature of the black hole. A positive specific heat indicates local stability, while a negative value implies instability. To examine global thermodynamic stability, one typically analyzes the Helmholtz free energy. However, as these features of RN-AdS black holes are well-established in the literature, we summarize the key outcomes: the SBH and LBH branches are locally thermodynamically stable due to their positive specific heat, whereas the IBH branch is unstable, as it is characterized by a negative specific heat.
In this work, we aim to investigate the potential connection between the thermodynamic phase transitions of black holes and the behavior of their associated quasi-periodic oscillations (QPOs). Our primary objective is to examine whether QPOs can serve as reliable indicators of black hole phase structure. To this end, we begin our study by evaluating the QPO frequencies within the Relativistic Precession (RP) models. Our analysis involves considering both massless and massive particles in orbit around the black hole. We extend our analysis to Warped Disk (WD), and Epicyclic Resonance (ER) frameworks also. The results are being presented concisely in the conclusion section.
II QPO and phase transition
Figure 1: Radial and vertical oscillatory response of the particle in a unstable circular null geodesics with different values of $Q$. Gray area represents the non-physical region where temperature is negative.
We begin by considering the Lagrangian for a test particle constrained to the equatorial plane, i.e., $\theta=\pi/2$. The Lagrangian is given by:
$\mathcal{L}=\frac{1}{2}\left[-f(r)\left(\frac{dt}{d\tau}\right)^{2}+\frac{1}{f(r)}\left(\frac{dr}{d\tau}\right)^{2}+r^{2}\left(\frac{d\phi}{d\tau}\right)^{2}\right],$ (8)
where $\tau$ denotes the proper time and $f(r)$ is the lapse function.
The canonical momenta associated with the generalized coordinates are computed as:
$p_{t}$ $=\frac{\partial\mathcal{L}}{\partial\dot{t}}=-f(r)\dot{t}=-E,$ (9)
$p_{r}$ $=\frac{\partial\mathcal{L}}{\partial\dot{r}}=\frac{1}{f(r)}\dot{r},$
$p_{\phi}$ $=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=r^{2}\dot{\phi}=L,$
where $E$ and $L$ represent the conserved energy and angular momentum of the particle, respectively. Here, a dot represents differentiation with respect to $\tau$.
Solving these equations yields the velocities:
$\dot{t}=\frac{E}{f(r)},\quad\dot{\phi}=\frac{L}{r^{2}}.$ (10)
We then proceed to compute the Hamiltonian using the Legendre transformation:
$2\mathcal{H}$ $=2(p_{t}\dot{t}+p_{r}\dot{r}+p_{\phi}\dot{\phi}-\mathcal{L})$ (11)
$=-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\phi}^{2}$
$=-\frac{E^{2}}{f(r)}+\frac{\dot{r}^{2}}{f(r)}+\frac{L^{2}}{r^{2}}=-\delta_{1},$
where we have used Eq. (10). Here, $\delta_{1}=1$ corresponds to timelike geodesics, and $\delta_{1}=0$ corresponds to null geodesics.
Introducing the effective potential $V_{r}$ for the radial motion through the relation $V_{r}=E^{2}-\dot{r}^{2}$, we obtain:
$V_{r}=f(r)\left[\delta_{1}+\frac{L^{2}}{r^{2}}\right].$ (12)
II.1 Massless particle (null geodesic) : RN black hole case
To study the QPO for massless particle we set $\delta_{1}=0$. The condition for unstable geodesic is $V^{\prime}_{r}(r_{0})=0$ and $V^{\prime\prime}_{r}(r_{0})<0$. From (12) we can have
$V^{\prime}_{r}(r_{0})=\frac{Lf^{\prime}(r_{0})}{r_{0}^{2}}-\frac{2Lf(r_{0})}{r_{0}^{3}}=0$ (13)
From Eq. (13) it is evident that $r_{0}$ is independent of L Using the lapse function of RN AdS black hole, Eq.(13) is solved for $r_{0}$ Now substituting the expression of mass in terms of event horizon radius $r_{+}$, we obtained the following relation between radius of the unstable circular orbit $r_{0}$ and event horizon radius $r_{+}$ as
$r_{0}=\frac{1}{2}\left(\sqrt{\frac{9\left(Q^{2}+r_{+}^{4}+r_{+}^{2}\right){}^{2}}{4r_{+}^{2}}-8Q^{2}}+\frac{3\left(Q^{2}+r_{+}^{4}+r_{+}^{2}\right)}{2r_{+}}\right)$ (14)
To analyze the fundamental frequencies associated with oscillatory motion of test particles around a RN AdS black hole, one considers small deviations from a unstable circular orbit. Specifically, the coordinates are perturbed as $r\to r_{0}+\delta r$ and $\theta\to\theta_{0}+\delta\theta$, where $r_{0}$ and $\theta_{0}$ define the equilibrium position. The effective potential $V_{\text{eff}}(r,\theta)$ can then be expanded in a Taylor series around this equilibrium configuration:
$V_{\text{eff}}(r,\theta)$ $=V_{\text{eff}}(r_{0},\theta_{0})+\delta r\left.\frac{\partial V_{\text{eff}}}{\partial r}\right|_{r_{0},\theta_{0}}+\delta\theta\left.\frac{\partial V_{\text{eff}}}{\partial\theta}\right|_{r_{0},\theta_{0}}$
$+\frac{1}{2}\delta r^{2}\left.\frac{\partial^{2}V_{\text{eff}}}{\partial r^{2}}\right|_{r_{0},\theta_{0}}+\frac{1}{2}\delta\theta^{2}\left.\frac{\partial^{2}V_{\text{eff}}}{\partial\theta^{2}}\right|_{r_{0},\theta_{0}}$
$+\delta r\delta\theta\left.\frac{\partial^{2}V_{\text{eff}}}{\partial r\partial\theta}\right|_{r_{0},\theta_{0}}+\mathcal{O}(\delta r^{3},\delta\theta^{3}).$ (15)
At the circular orbit, the first-order derivatives of the potential vanish due to equilibrium conditions, and the higher-order terms are neglected under the small perturbation assumption. As a result, the equations of motion for the radial and vertical perturbations reduce to those of a harmonic oscillator. These perturbations manifest as oscillations with characteristic frequencies as measured by a distant observer [77]:
$\frac{d^{2}\delta r}{dt^{2}}+\Omega_{r}^{2}\delta r=0,\quad\frac{d^{2}\delta\theta}{dt^{2}}+\Omega_{\theta}^{2}\delta\theta=0,$ (16)
where the squared radial and vertical frequencies are given by [76]
$\Omega_{r}^{2}=-\frac{1}{2g_{rr}\dot{t}^{2}}\left.\frac{\partial^{2}V_{r}}{\partial r^{2}}\right|_{r_{0},\theta=\pi/2},$ (17)
$\Omega_{\theta}^{2}=-\frac{1}{2g_{\theta\theta}\dot{t}^{2}}\left.\frac{\partial^{2}V_{r}}{\partial\theta^{2}}\right|_{r_{0},\theta=\pi/2}.$ (18)
$\Omega_{\phi}^{2}=\frac{-\partial_{r}g_{tt}}{\partial_{r}g_{\phi\phi}}=\sqrt{\frac{f^{\prime}(r)}{2r}}.$ (19)
Here $\Omega_{\phi}=\Omega_{\theta}=\Omega_{K}$. $\Omega_{K}$ is the Keplerian frequency which is the angular velocity of a test particle revolving around a black hole, as observed by a distant observer.
Figure 2: Behavior of the quasi-periodic oscillation (QPO) frequencies of both upper $\nu_{U}$ and lower $\nu_{L}$ for RN AdS black holes as functions of temperature, across different black hole branches: small (SBH), intermediate(IBH) and large (LBH). Here, we have considered $Q=0.05$.
These expressions characterize the radial and vertical oscillatory response of the particle to small disturbances from the circular orbit in the equatorial plane. To convert these frequencies into physical frequencies in units of Hertz (Hz), we employ the following relation:
$\nu_{i}=\frac{c^{3}}{2\pi GM}\cdot\Omega_{i}$ (20)
Next we calculate the temperature of the RN black hole using the following formula :
$T=\frac{-Q^{2}+3r_{+}^{4}+r_{+}^{2}}{4\pi r_{+}^{3}}$ (21)
it is evident that at $r_{+}=\frac{\sqrt{\sqrt{12Q^{2}+1}-1}}{\sqrt{6}}$, temperature goes to zero. Hence $r_{+}<\frac{\sqrt{\sqrt{12Q^{2}+1}-1}}{\sqrt{6}}$ is the unphysical region.Now we plot the radial and vertical oscillatory response of the particle in a unstable circular null geodesics with different values of $Q$ in Fig.1. The distinction between physical area and non physical region is clearly observable in the figure.
II.2 QPO Models
In this section, we explore the behavior of twin-peak quasi-periodic oscillations (QPOs) in the background of Reissner–Nordström (RN) black holes. The upper ($\nu_{U}$) and lower ($\nu_{L}$) QPO frequencies are modeled as functions of the radial coordinate and black hole parameters, guided by several well-established theoretical frameworks [78].
The QPO models considered in our analysis are summarized as follows [78]:
  • Relativistic Precession (RP) model: In this framework, the upper and lower frequencies correspond to the azimuthal and periastron precession frequencies respectively, i.e., $\nu_{U}=\nu_{\phi}$, $\nu_{L}=\nu_{\phi}-\nu_{r}$.
  • Epicyclic Resonance (ER) models: These models describe resonant interactions within a geometrically thick disk. The different variants are:
    • ER2: $\nu_{U}=2\nu_{\theta}-\nu_{r}$, $\nu_{L}=\nu_{r}$,
    • ER3: $\nu_{U}=\nu_{\theta}+\nu_{r}$, $\nu_{L}=\nu_{\theta}$,
    • ER4: $\nu_{U}=\nu_{\theta}+\nu_{r}$, $\nu_{L}=\nu_{\theta}-\nu_{r}$.
  • Warped Disk (WD) model: This model, based on motion in a thin accretion disk, defines the frequencies as $\nu_{U}=2\nu_{\phi}-\nu_{r}$ and $\nu_{L}=2(\nu_{\phi}-\nu_{r})$.
The primary goal of this study is to determine the behavior of the upper and lower QPO frequencies across these different models and examine how they respond to changes in black hole’s thermodynamic phases. By comparing the evolution of $\nu_{U}$ and $\nu_{L}$ across different black hole branches, we aim to identify possible imprints of phase transitions in the QPO spectrum. In doing so, we assess whether these oscillation models can provide insight into the thermodynamic stability and structure of the RN black hole phases.
II.3 Signatures of Phase transition
We begin our analysis with the Relativistic Precession (RP) model, where the upper and lower quasi-periodic oscillation (QPO) frequencies are defined as:
$\nu_{U}=\nu_{\phi},\qquad\nu_{L}=\nu_{\phi}-\nu_{r}.$
By plotting these frequencies against the temperature for subcritical charge values (e.g., $Q=0.05$), a clear separation of black hole phases becomes evident within the frequency domain.
For $Q=0.05$, the phase structure can be delineated as follows: the small black hole (SBH) branch exists from $(r_{+},T)=(0.0498149,0)$ to $(0.08761,0.633382)$, the intermediate black hole (IBH) branch spans from $(0.08761,0.633382)$ to $(0.57066,0.274613)$, and the large black hole (LBH) branch extends from $(0.57066,0.274613)$ towards the asymptotic limit $(r_{+},T)\rightarrow(\infty,\infty)$.
As illustrated in Fig. 4, the solid red curve represents the SBH branch. In this regime, both $\nu_{U}$ and $\nu_{L}$ decrease with rising temperature, indicating that the frequency of a particle orbiting the black hole reduces as the system heats up. This monotonic decrease is a hallmark of stability in the small black hole phase.
In contrast, along the IBH branch(solid blue curve), both frequencies increase with temperature, suggesting an unstable regime where the particle experiences stronger oscillations — a behavior consistent with thermodynamic instability. Finally, in the LBH branch (solid black curve) the frequencies begin to decrease once more and eventually saturate, reflecting a return to stability in the large black hole phase. When the black hole charge exceeds its critical value, the system exhibits a single thermodynamically stable phase across the entire temperature domain. This is reflected in the frequency behavior as well: for $Q$ above the critical threshold, both $\nu_{U}$ and $\nu_{L}$ decrease monotonically with temperature and asymptotically approach constant values. This smooth decline further confirms the absence of phase transitions and supports the thermodynamic stability of the system in this regime.
This analysis highlights that the QPO behavior of a test particle can serve as an effective probe of black hole phase transitions and their stability. The connection between dynamical frequencies and thermodynamic stability suggests a compelling observational signature for identifying black hole phases. However, further observational data will be crucial to solidify this correspondence.
To reinforce our findings, we provide a numerical table below summarizing the QPO frequencies across the three branches, which further illustrates the thermodynamic phase structure inherent in RN AdS black holes.
Black Hole Phase Temperature ($T$) Upper Frequency ($\nu_{U}$) [Hz] Lower Frequency ($\nu_{L}$) [Hz]
$Q<Q_{c}$ (Three-Phase Structure)
0.0119366 164709 47810.3
0.41958 159488 41240.9
SBH 0.573522 149007 30901.6
0.625255 137596 22577.1
0.633382 $(T_{C})$ 129336 17886.8
0.632781 126880 16665.2
0.620704 117308 12539.0
IBH 0.420766 66966.8 1710.48
0.329510 49723.0 514.412
0.274613 $(T_{C})$ 36308.2 70.2909
0.274948 35784.1 59.5509
0.318111 32972.8 9.14767
LBH 0.742716 32382.3 0.0399761
1.209580 32379.8 0.00212883
1.682490 32379.6 0.000293692
$Q>Q_{c}$ (Single Stable Phase)
0.119366 34981.8 5588.86
0.156414 34701.6 4759.03
Stable BH 0.183765 34399.5 3948.53
0.205161 34108.0 3235.86
0.222794 33843.0 2639.13
Table 1: Numerical values of upper and lower QPO frequencies of massless particle across different black hole phases for $Q<Q_{c}$ (SBH, IBH, LBH) and $Q>Q_{c}$ (single-phase stable branch). The trends in QPO behavior reflect the underlying thermodynamic phase structure and stability patterns of RN-AdS black holes.
We have also performed our analysis using the ER2, ER3, ER4 models as well as the Warped Disk (WD) model. All these QPO models exhibit consistent qualitative behavior and successfully capture the features of black hole phase transitions and the associated stability structure. For the sake of brevity, we omit the detailed results here. It is worth noting, however, that while the overall trends remain coherent across models, the actual frequency values may vary. Among them, the ER2 model yields the highest frequency values. Across all five models, the behavior of the upper and lower QPO frequencies as functions of temperature remains qualitatively similar: in the SBH phase, frequencies decrease with temperature, indicating stable motion; in the IBH phase, the frequencies rise, signaling instability; and in the LBH phase, they gradually level off and approach constant values in the asymptotic limit, once again reflecting stability.
III Massive particle(timelike geodesic)
Figure 3: Radial and vertical oscillatory response of the particle in a unstable circular timelike geodesics with different values of $Q$.We have considered L=20. Gray area represents the non-physical region where temperature is negative.
In this section, we perform similar analysis for massive particle where we set $\delta_{1}=0$ in Eq.(12) The expressions for E and L is obtained to be :
$E^{2}=\frac{2f\left(r_{0}\right){}^{2}}{2f\left(r_{0}\right)-r_{0}f^{\prime}\left(r_{0}\right)},$ (22)
and
$L^{2}=\frac{r_{0}^{3}f^{\prime}\left(r_{0}\right)}{2f\left(r_{0}\right)-r_{0}f^{\prime}\left(r_{0}\right)}.$ (23)
Next, using the expression of L, the relation between $r_{0}$ and $r_{+}$ is obtained.Here the $r_{0}$ depends upon the angular momentum $L$ unlike massless particle case. Using the obtained relation between event horizon radius and radius of the unstable circular orbit we obtained the expressions for radial and vertical frequencies as shown in the previous section.In Fig.3, we plot the frequencies as a function of event horizon radius $r_{+}$. The black line represents the temperature zero condition and the shaded portion is the non physical region.
Figure 4: Behavior of the quasi-periodic oscillation (QPO) frequencies of both upper $\nu_{U}$ and lower $\nu_{L}$ for RN black holes as functions of temperature, across different black hole branches: small (SBH), intermediate(IBH) and large (LBH). Here, we have considered $Q=0.05$.
Now we compute the upper and lower QPO frequencies within the Relativistic Precession (RP) model for massive particles, where the identification of frequencies remains the same:
$\nu_{U}=\nu_{\phi},\qquad\nu_{L}=\nu_{\phi}-\nu_{r}.$
As in the massless case, we analyze the behavior of these frequencies as functions of the black hole temperature for a subcritical charge value, e.g., $Q=0.05$. The frequency plots once again reveal the underlying phase structure of the black hole: the small black hole (SBH) branch spans from $(r_{+},T)=(0.0498149,0)$ to $(0.08761,0.633382)$, followed by the intermediate black hole (IBH) branch from $(0.08761,0.633382)$ to $(0.57066,0.274613)$, and finally the large black hole (LBH) branch extends from $(0.57066,0.274613)$ toward the asymptotic regime.
In the SBH region, both $\nu_{U}$ and $\nu_{L}$ decrease with increasing temperature, indicating dynamical stability. In the IBH branch, the frequencies increase with temperature, which corresponds to an unstable regime where orbital motion becomes more sensitive to perturbations. Interestingly, as the system transitions into the LBH branch, the lower frequency $\nu_{L}$ for massive particles drops sharply and reaches zero exactly at the onset of the LBH phase. This vanishing of $\nu_{L}$ signifies that the particle enters a state of ultimate stability, where radial oscillations effectively cease. Beyond this point, $\nu_{L}$ becomes negative, which is unphysical and indicates the breakdown of the oscillatory interpretation in this regime. In contrast, the upper frequency $\nu_{U}$ continues to follow the same trend as in the massless case, decreasing monotonically and asymptotically approaching a constant value.
For charge values above the critical threshold ($Q>Q_{c}$), the temperature-frequency relation is smooth and monotonic, with no indication of a phase transition. In this regime, both frequencies evolve predictably: $\nu_{U}$ decreases gradually with temperature, while $\nu_{L}$ remains positive and asymptotically approaches zero from above, indicating a stable configuration throughout.
In summary, while the overall qualitative behavior of QPO frequencies for massive particles is consistent with the massless case, the vanishing of $\nu_{L}$ at the LBH threshold introduces a distinctive physical signature. A numerical summary of the results for massive particles is provided in the table below.
Black Hole Phases Temperature (T) Lower Frequency ($\nu_{L}$) [Hz] Upper Frequency ($\nu_{U}$) [Hz]
0.0119366 47810.1 164711.
0.41958 41240.4 159490.
SBH 0.573522 30900.7 149009.
0.625255 22575.7 137599.
0.633382($T_{C}$) 17885.1 129339.
0.632781 16663.4 126883.
0.620704 12536.7 117311.
IBH 0.420766 1701.11 66975.1
0.32951 486.214 49738.1
0.280928 0 39376.7
0.274613($T_{C}$) -338.359 36366.1
$Q>Q_{c}$ (Single Stable Phase)
0.119366 4966.37 35053.9
0.156414 3946.51 34783.2
Stable Phase 0.183765 2845.78 34494.1
0.205161 1715.0 34218.4
0.222794 537.2 33972.2
Table 2: Numerical values of the temperature of RN AdS black hole and QPO frequencies for massive particle
We have also extended our analysis to other QPO models, including the Warped Disk (WD) model and the Epicyclic Resonance (ER) models. Among them, the WD model exhibits behavior that closely mirrors that of the Relativistic Precession (RP) model. Both models consistently capture the thermodynamic stability pattern of black hole phases and support our interpretations based on frequency evolution. In contrast, the ER modelsnamely ER2, ER3, and ER4 do not fully align with the expected behavior. Specifically, in all ER models, the frequencies associated with the large black hole (LBH) phase exhibit an increasing trend with temperature, which contradicts the physical interpretation drawn from the RP and WD models. While the presence of distinct phase structures remains evident across all models, the discrepancies observed in the ER frameworks raise an interesting question that needs to be addressed in our future works.
IV Summary and Concluding Remarks
In this work, we probe the well-known thermodynamic phase structure of the RN-AdS black hole through the lens of its quasi-periodic oscillations (QPOs). Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational pattern? These were the central questions guiding our study. Beginning with the Relativistic Precession (RP) model, we demonstrated that the behavior of both upper and lower QPO frequencies, when plotted against temperature reveals a clear imprint of the underlying phase structure for subcritical charge values. Specifically, the frequencies trace out the small, intermediate, and large black hole branches in a manner that reflects their respective thermodynamic stability. In the small and large black hole phases, the frequencies decrease with temperature, indicating stability, whereas the intermediate branch is characterized by increasing frequencies, signaling instability. For charge values above the critical point, our results show a smooth, monotonic behavior of the frequencies with no evidence of phase transitions, further supporting the interpretation of a single thermodynamically stable phase in this regime. We have also extended our analysis to massive particles and confirmed that the qualitative features of the frequency-temperature relationship remain intact. In addition to the RP model, we examined the behavior of QPO frequencies under the Warped Disk (WD) and Epicyclic Resonance (ER) models. For massless particles, all models including ER2, ER3, ER4, and WD exhibit consistent qualitative trends. In the massive particle case, however, the ER models exhibit noticeable deviations most prominently in the LBH branch, where the QPO frequencies increase with temperature, in sharp contrast to the stability characteristics captured by the RP and WD models. On the other hand, the WD model continues to mirror the RP model and successfully reflects the expected stability pattern. Although the presence of distinct phase structures remains evident across all models which is an encouraging feature for further exploration in other black hole systems. The precise relationship between thermodynamic stability and QPO behavior may vary depending on the black hole geometry or the choice of thermodynamic ensemble. This remains an open question, inviting deeper investigation in future studies.
These findings suggest that QPO frequencies may serve as powerful observational tools for probing the thermodynamic phase structure and stability of black holes. With the aid of available observational data, the imprints of phase transitions can be further substantiated, and the size of different black hole phases may be constrained. This presents a promising avenue for future research and warrants more detailed investigation.
Acknowledgements
BH would like to thank DST-INSPIRE, Ministry of Science and Technology fellowship program, Govt. of India for awarding the DST/INSPIRE Fellowship[IF220255] for financial support.